Calculate the equilibrium concentrations of H2, I2, and HI at 700 K if the initial concentrations are [H2] = 0.200 M and [I2] = 0.400 M. The equilibrium constant Kc for the reaction following reaction is 57.0 at 700 K. (Show Work)
H2(g)+I2(g)<---
                   ---->2HI(g)
Answer #1

We have,

Kc = [HI]^2/[H2][I2] = 57

ICE chart

              H2 +   I2   <===> 2HI

I             0.2      0.4                -

C            -x        -x               -2x

E          0.2-x   0.4-x             2x

So,

57 = (2x)^2/(0.2-x)(0.4-x)

57x^2 - 34.2x + 4.56 = 4x^2

53x^2 - 34.2x + 4.56 = 0

x = 0.19 M

equilibrium concentration of,

[H2] = 0.2 - 0.19 = 0.01 M

[I2] = 0.4 - 0.19 = 0.21 M

[HI] = 2 x 0.19 = 0.38 M